# Scalable Training of Trustworthy and Energy-Efficient Predictive Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

### v3

@@ III-B Data Cleaning and Pre-processing
Total 12,421
TABLE II: Number of data samples discarded from each dataset due to values of the $L2$-norm (also known as spectral norm) of the force tensor being unreasonably over 100 eV/angstrom.
The histogram of the values of the energy per atom and $L^{2}$-norm of forces, after removing data samples with force values unreasonably high are shown in Figure 3. The partition of each dataset into training, validation, and testing has been performed using the 80%-10%-10% splitting. In Figure 4 we show the histograms of the distribution of energies and atomic forces for the training, validation, and testing subsets of each dataset.
Figure 3: Normalized histograms for each dataset of the total energy per atom (top) and the $L^{2}$-norm of the force tensor (bottom) for each dataset after removing data samples with $L^{2}$-norm of the force tensor unreasonably higher than 100 eV/angstrom.
@@
III-C Scalable Data Management
@@
III-C1 ADIOS for High Performance I/O
Several publicly available atomistic materials modeling datasets are stored using bespoke schemas and exported as large collections of files. For example, the OC2020 dataset [43] consists of over 50,000 files. Storing multiple such datasets adds prohibitively high metadata overhead on the PFS and leads to slow data ingestion during the training process. For efficiently storing and reading of large training data, we use the ADIOS  [69] scientific data management library, which provides a state-of-the-art solution for managing extreme-scale data. ADIOS is designed to provide scalable I/O on the largest supercomputers in the world and has been successfully used in science applications that write and read several petabytes in a single simulation run.
@@ III-C1 ADIOS for High Performance I/O
The number of sub-files controls the concurrency level while reading data in parallel. This $n:m$ pattern in which $n$ processes concurrently read data from $m$ sub-files is pivotal to obtaining high reading performance using ADIOS, which provides several options to tune I/O performance, including configuring the number of sub-files. We create the graph structures from input data and store them in ADIOS as a separate pre-processing step. We have developed a data writer and reader in HydraGNN for writing and reading graph data, respectively, from ADIOS files during the training process. When an ADIOS file is created, we split the data samples into three groups - ‘trainset’ representing training data, ‘valset’ for data used for validation, and ‘testset’ data for testing the model performance. This logical grouping of data samples helps us read different groups of data samples for different tasks during the training process.
III-C2 DDStore
@@
Figure 5: Different approaches for shuffling data during the training process. In a), data is read from the shared file system in which each graph object is stored in its own separate file (high file system metadata overhead). In b), data is read from an ADIOS file (low metadata overhead, high I/O bandwidth). In c) all data is read once into DDStore [70], an in-memory data store which uses MPI one-sided RMA operations to obtain data from remote processes (best performance).
@@
DDP [71, 72, 73, 74, 75] involves distributing training data amongst the available compute resources. Data is grouped into batches, and GPUs train on one batch at a time before fetching the next batch until all batches are processed in an epoch. Frequently reading data from the PFS, even via a high-performance library such as ADIOS, is an expensive operation because I/O over the shared PFS is the slowest operation in a computing system.
@@
To provide fast data retrieval during training, we used DDStore [70], a distributed data store that provides in-memory data transfer between processes. When training begins, processes read data from ADIOS files and load into the node’s memory, which maintains a global map of data samples on each process. When a GPU requests a new batch of data, DDStore fetches the data from remote processes using low latency, fast communication techniques instead of reading data from the PFS. By restricting access to the PFS to the initial bootup phase, DDStore ensures that obtaining a batch is a fast, in-memory operation. Experiments described in [70] show that it leads to a 6$\times$ speedup in overall training time.
@@
DDStore provides options to tune the size of data chunks stored on each process (chunking), replicating a dataset on internal sub-groups of processes (replication), and the communication mechanism selected for fetching data. For our experiments, data is split evenly amongst all processes, and a single replica of the dataset is maintained across all processes. For efficient data retrieval, the low latency MPI one-sided remote memory access (RMA) operations were used. Fig. 5 shows the data loading and caching approach used by DDStore compared to traditional approaches that read data directly from the file system. Section IV shows the time taken to obtain a batch of data samples for different model sizes and node counts.
III-D MTL for simultaneous prediction of energies and atomic forces
@@
In HydraGNN, each predicted quantity is associated with a separate loss function and the global objective function minimized during the training is a linear combination of the individual loss functions. Denoting with $\mathbf{F}\in\mathbb{R}^{n\times 3}$ the atomic force tensor what contains atomic forces acting on all the $N$ atoms of a structure
@@
where $T$ represents the transposition of a vector. Denoting with $N$ the total number of parameters to train in the GFM architecture, the MTL loss function $\ell_{\mathrm{MTL}}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{+}$ is:
$\ell_{\mathrm{MTL}}(\mathbf{W})=\alpha_{\mathrm{energy}}\lVert e_{\textrm{predict}}(\mathbf{W})-e\rVert_{1}+$ (3)
$\alpha_{\mathrm{forces}}\lVert\mathbf{F}_{\textrm{predict}}(\mathbf{W})-\mathbf{F}\rVert_{1},$ (4)
where $\lVert\cdot\rVert_{1}$ represents the $L_{1}$ norm of a vector or the induced tensor norm, $e$ and $\mathbf{F}$ represent the true values of energies and forces, $\mathbf{e}_{\textrm{predict}}$ and $\mathbf{F}_{\textrm{predict}}$ are the corresponding predictions given by HydraGNN, and the scalar positive weights $\alpha_{\mathrm{energy}}$ and $\alpha_{\mathrm{forces}}$ are used to calibrate the individual terms of the loss function for each individual properties. Similarly to [77], we calibrate the scalar positive weights $\alpha_{\mathrm{energy}}$ and $\alpha_{\mathrm{forces}}$ to account for the fact that the energy is a global quantity, while the atomic forces are local quantities. More specifically, we set the values of these scalar weights to $\alpha_{\mathrm{energy}}=1$ and $\alpha_{\mathrm{forces}}=100$. While it is helpful to optimize the weights as a hyperparameter, we found that this set of values is a suitable choice. An illustration of the main components of the HydraGNN architecture used for MTL is described in Figure 6.
@@ III-E Scalable HPO
To perform HPO at large scale, we used DeepHyper [84], an open-source Python package designed for optimizing hyperparameters, searching for optimal neural architectures, and quantifying uncertainty through the deep ensembles. Specifically, we used asynchronous Bayesian optimization that continuously refines a surrogate model by sampling hyperparameter configurations. The efficacy of DeepHyper’s asynchronous Bayesian optimization has been demonstrated across various deep learning benchmarks, outperforming methods such as random search, genetic algorithms, and Hyperband in environments equipped with CPUs and GPUs. In the DeepHyper setup, a manager node refines the surrogate model and suggests promising configurations while worker nodes perform the evaluations. Our approach uses a centralized architecture with process-based parallelism, optimizing the allocation of tasks across computing nodes to avoid bottlenecks.
@@
III-F Scalable UQ with GNN Ensembles
Ensemble methods are widely utilized in UQ to compile predictions from various models, termed ensemble members, into a unified forecast. The goal of these methods is to enhance model generalization by drawing on the diverse capabilities of each individual model [88]. To promote a varied set of predictions, practices such as different model initializations, techniques like Bagging and Boosting, and the integration of diverse network architectures are employed. Research conducted by Egele et al. [89] demonstrated that expanding the variety of network architectures within an ensemble can improve the diversity, thereby increasing the precision of uncertainty assessments. They also developed a technique for concurrently training multiple candidate models, which optimizes the use of computational resources. Ensemble methods are acknowledged for their ability to deliver reliable uncertainty estimates and their ease of implementation and scalability, making them practical for various UQ applications.
@@ III-F Scalable UQ with GNN Ensembles
and the uncertainty is measured as the standard deviation,
@@
IV Performance Measurements
IV-A Setup
For strong and weak scaling studies we utilize three different sizes of GFM architectures, denoted as SMALL, MEDIUM, and LARGE. They differ in the total number of parameters, ranging from approximately 60,000 to 163 million. Table III provides details about the three model sizes.
Experiments were conducted on two DOE supercomputers: Frontier at ORNL and Perlmutter at NERSC. Both systems provide state-of-the-art GPU-based heterogeneous architectures. Frontier, located at the Oak Ridge Leadership Computing Facility at ORNL, is currently the world’s fastest and most powerful supercomputer [90]. It comprises a total of 9,408 compute nodes, each featuring a single 64-core AMD EPYC 7763 (Milan) CPU and four AMD Instinct MI250X GPU accelerators, effectively providing eight GPU units per node. Running with one rank per GPU unit, each rank has 64 GB of DDR4 (CPU) and 64 GB of HBM2e (GPU) memory.
@@ IV-A Setup
To aid in monitoring HydraGNN execution in real-time for a subset of the analysis carried out on Frontier, an AMD Research utility, Omnistat [91], was employed to sample a variety of GPU telemetry metrics including occupancy, high-bandwidth memory (HBM) usage, power, temperature, energy consumption, and clock/memory frequencies on a per GCD basis across all nodes assigned to an individual run. This Python-based utility was executed entirely in user-space implemented as a Prometheus client on each assigned compute node and combines low-overhead sampling via AMD’s system management interface (SMI) at fixed intervals with a temporary Prometheus server [92] instantiated on one CPU core of the master compute host per batch job. Minimal job overhead (less than 0.5%) was observed when running HydraGNN training with this approach for sampling intervals down to one second.
@@
TABLE III: GNN model sizes used for strong and weak scaling tests on NERSC-Perlmutter and OLCF-Frontier.
IV-B I/O Performance for Reading Large Data
In Section III-C, we described using the ADIOS scientific data management library for fast storage and retrieval of large training data. In this section, we show the performance of reading large data in HydraGNN for training models.
Of all the datasets used in this study, the Open Catalyst 2020 dataset is the largest in terms of the number of atomistic structures, the storage size of the dataset, and the number of files across which data is stored. The original dataset consists of over 50,000 files. The dataset was pre-processed into ADIOS and was configured to use just over 50 ADIOS sub-files, which led to a 1000$\times$ reduction in the metadata footprint.
@@ IV-B I/O Performance for Reading Large Data
IV-C HydraGNN Training Scaling Results
We now analyze the scaling performance of HydraGNN on Frontier and Perlmutter. We present weak and strong scaling trends, along with a breakdown of component operations in HydraGNN as we scale it up. Experiments were performed with up to 2,048 nodes on Frontier and 256 nodes on Perlmutter using the three model sizes discussed in Table III.
IV-C1 Weak Scaling
For the weak scaling runs, we configured each GPU to process 3,500 atomistic structures equally. Fig. 9 shows the weak scaling performance on Perlmutter and Frontier as we vary the number of GPUs used for the training. The reported time represents the average training time per epoch. We conducted experiments with up to 2,048 GPUs on Frontier and 1,024 GPUs on Perlmutter. The number of GPUs on Perlmutter was due to constraints on available node hours. We observe that the parallel efficiency of weak scaling experiments drops as we increase the number of GPUs beyond 256 for both Perlmutter and Frontier. This is attributed to increased communication costs as we scale the number of GPUs and the overhead associated with using varying graph sizes.
Figure 9: Weak scaling of HydraGNN multitasking pretraining on a problem of size 3,500 atomistic structures per GPU for (top) Frontier and (bottom) Perlmutter.
@@ IV-C2 Strong Scaling
Figure 11: Strong scaling of HydraGNN multitasking pretraining on a problem of 120 million graphs on Frontier and 2 million graphs on Perlmutter with three GNN model sizes.
The drop in scaling performance is attributed to load imbalance - an artifact of small graph sizes. As shown in Fig. 1, we use a diverse dataset where graph sizes vary by up to 400 nodes, and the number of edges in the larger graphs exceeds 12,500. This results in an imbalanced workload among GPUs in each batch, causing some GPUs to finish training before others. As GPUs must synchronize for exchanging model weights, the runtime is dominated by the GPUs that must train on larger graphs. Effectively, this leads to sub-standard utilization of compute resources and poses a challenge towards achieving high-performant, scalable training. Training on large volumes of data can help develop robust models because of the diverse nature of data, but it affects the computational performance as the workload can vary greatly.
@@
The EGNN model we used is particularly vulnerable to this problem. The time required for forward calculations in EGNN is directly proportional to the number of edges in the graph sample. For datasets with highly variable edge counts between graph samples, the likelihood of load imbalance between GPUs increases. Fig. 12 illustrates the time spent on the forward task in the EGNN model with different model sizes. We observe an almost linear relationship between forward time and graph size at each batch (measured by the number of edges). The SMALL models show large variances on both machines, which is expected due to system noise being more consequential with smaller model sizes. Significant performance differences (e.g., the difference between minimum and maximum time) are observed due to the varying graph sizes in our datasets. However, for other tasks (data loading and backward), we do not observe a similar correlation, as they are agnostic of the graph size. Fig. 13 illustrates the average percentage of time spent waiting during three tasks: data loading, forward pass, and backward pass. It highlights a significant waiting period during the forward pass, primarily due to varying graph sizes. This waiting time increases as the disparity in graph sizes among GPUs grows. Other tasks, such as data loading and backward pass, also involve waiting time, but to a much lesser extent.
@@
where $T_{max}$ and $T_{avg}$ represent the maximum runtime and the average runtime for training an epoch, respectively, among all computing resources (GPUs in our case). These timings represent the time to perform training (forward and backward calculations) and do not include wait times during synchronization. For a well-balanced workload, LIF approaches 1.0 from above, whereas it increases as the workload imbalance increases. Fig. 14 presents the LIF scores that show the imbalance among processes. The trend remains consistent: while data loading and backward pass exhibit nearly balanced workloads (with scores close to 1.0), the forward pass shows imbalanced workload characteristics as it deviates from 1.0.

### v4

@@ 3.2 Data cleaning and pre-processing
Total 12,422
Table 2: Number of data samples discarded from each dataset due to values of the $L2$-norm (also known as spectral norm) of the force tensor being unreasonably over 100 eV/angstrom.
Datasets from different sources were generated with different electronic structure methods, leading to global shifts in the energy calculated for each element of the periodic table, and consequently also on the energies of the organic and inorganic compounds obtained by combining multiple elements. In order to re-align the multi-source multi-fidelity data, we adopted the procedure proposed in [65] to transform the energy of each atomistic structure during pre-processing by subtracting a linear regression term. For each dataset, the linear regression term was calculated by solving the following least-squares problem:
$\mathop{\mathrm{arg\,min}}_{C_{1},\ldots,C_{118}}\sum_{i=1}^{N_{data}}\bigg{(}e^{i}_{0}-\sum_{Z=1}^{118}C_{Z}n^{i}_{Z}\bigg{)}^{2}$ (1)
where $e^{i}_{0}$ is the reference energy for the atomistic structure $i$, and $n^{i}_{Z}$ counts the number of atoms of element number $Z$ belonging to the atomistic structure $i$. By separately determining the regression coefficients $C_{Z}$ (with the dimension unit of eV/atom) for each dataset, we were able to remove the largest source of variability between datasets. This led to a more consistent training of the GFMs using the labeled energy values. Even if the datasets cover two-thirds of the natural elements of the periodic table, we built the linear regression model to account for all the 118 elements (natural and artificial) of the periodic table to ensure that our software infrastructures accommodates future generalizations of our current approach.
The histogram of the values of the energy per atom before and after removing the linear regression term and $L^{2}$-norm of forces, after removing data samples with force values unreasonably high are shown in Figure 3. The partition of each dataset into training, validation, and testing has been performed using the 80%-10%-10% splitting. In Figure 4 we show the histograms of the distribution of energies and atomic forces for the training, validation, and testing subsets of each dataset.
Figure 3: Normalized histograms for each dataset of the total energy per atom (top) , energy per atom after removing the linear regression term from each dataset (center), and the $L^{2}$-norm of the force tensor (bottom) for each dataset after removing data samples with $L^{2}$-norm of the force tensor unreasonably higher than 100 eV/angstrom.
@@
3.3 Scalable data management
@@
3.3.1 ADIOS for high performance I/O
Several publicly available atomistic materials modeling datasets are stored using bespoke schemas and exported as large collections of files. For example, the OC2020 dataset [39] consists of over 50,000 files. Storing multiple such datasets adds prohibitively high metadata overhead on the PFS and leads to slow data ingestion during the training process. For efficiently storing and reading large volumes of training data, we use the ADIOS  [66] scientific data management library, which provides a state-of-the-art solution for managing extreme-scale data. ADIOS is designed to provide scalable I/O on the largest supercomputers in the world and has been successfully used in science applications that write and read several petabytes in a single simulation run.
@@ 3.3.1 ADIOS for high performance I/O
The number of sub-files controls the concurrency level while reading data in parallel. This $n:m$ pattern in which $n$ processes concurrently read data from $m$ sub-files is pivotal to obtaining high reading performance using ADIOS, which provides several options to tune I/O performance, including configuring the number of sub-files. We create the graph structures from input data and store them in ADIOS as a separate pre-processing step. We have developed a data writer and reader in HydraGNN for writing and reading graph data, respectively, from ADIOS files during the training process. When an ADIOS file is created during the data pre-processing, we split the data samples into three groups - ‘trainset’ representing training data, ‘valset’ for data used for validation, and ‘testset’ data for testing the model performance. This logical grouping of data samples helps us read different groups of data samples for different tasks during the training process.
3.3.2 DDStore
@@
Figure 5: Different approaches for shuffling data during the training process. In a), data is read from the shared file system in which each graph object is stored in its own separate file (high file system metadata overhead). In b), data is read from an ADIOS file (low metadata overhead, high I/O bandwidth). In c), all data is read once into DDStore [67], an in-memory data store which uses MPI one-sided RMA operations to obtain data from remote processes (best performance).
@@
DDP [68, 69, 70, 71, 72] involves distributing training data amongst the available compute resources. Data is grouped into batches, and GPUs train on one batch at a time before fetching the next batch until all batches are processed in an epoch. Frequently reading data from the PFS, even via a high-performance library such as ADIOS, is an expensive operation because I/O over the shared PFS is one of the slowest operations in a computing system.
@@
To provide fast data retrieval during training, we used DDStore [67], a distributed data store that provides in-memory data transfer between processes. When training begins, processes read data from ADIOS files and load into the node’s memory, which maintains a global map of data samples on each process. When a GPU requests a new batch of data, DDStore fetches the data from remote processes using low latency, fast communication techniques rather than reading data from the PFS. By restricting access to the PFS to the initial bootup phase, DDStore ensures that obtaining a batch is a fast, in-memory operation. Experiments described in [67] show that it leads to a 6$\times$ speedup in overall training time.
@@
DDStore provides options to tune the size of data chunks stored on each process (chunking), replicating a dataset on internal sub-groups of processes (replication), and the communication mechanism selected for fetching data. For our experiments, data is split evenly amongst all processes, and a single replica of the dataset is maintained across all processes. For efficient data retrieval, the low latency MPI one-sided remote memory access (RMA) operations were used. Fig. 5 shows the data loading and caching approach used by DDStore compared to traditional approaches that read data directly from the file system. Section 4 shows the time taken to obtain a batch of data samples for different model sizes and node counts.
3.4 MTL for prediction of energy and atomic forces
@@
In HydraGNN, each predicted quantity is associated with a separate loss function and the global objective function minimized during the training is a linear combination of the individual loss functions. We denote by $\mathbf{F}\in\mathbb{R}^{n\times 3}$ the atomic force tensor that contains the atomic forces acting on all the $N$ atoms
@@
where $T$ represents the transposition of a vector. Denoting by $P$ the total number of parameters to train in the GFM architecture, the MTL loss function $\ell_{\mathrm{MTL}}:\mathbb{R}^{P}\rightarrow\mathbb{R}^{+}$ is:
$\ell_{\mathrm{MTL}}(\mathbf{W})=\alpha_{\mathrm{energy}}\lVert e_{\textrm{predict}}(\mathbf{W})-e\rVert_{1}+\alpha_{\mathrm{forces}}\lVert\mathbf{F}_{\textrm{predict}}(\mathbf{W})-\mathbf{F}\rVert_{1},$ (4)
where $\lVert\cdot\rVert_{1}$ represents the $L_{1}$ norm of a vector or the induced tensor norm, $e$ and $\mathbf{F}$ represent the true values of energies and forces, $\mathbf{e}_{\textrm{predict}}$ and $\mathbf{F}_{\textrm{predict}}$ are the corresponding predictions given by HydraGNN, and the scalar positive weights $\alpha_{\mathrm{energy}}$ and $\alpha_{\mathrm{forces}}$ are used to calibrate the individual terms of the loss function for each individual properties. Similarly to [65], we calibrate the scalar positive weights $\alpha_{\mathrm{energy}}$ and $\alpha_{\mathrm{forces}}$ to account for the fact that the energy is a global quantity, while the atomic forces are local quantities. More specifically, we set the values of these scalar weights to $\alpha_{\mathrm{energy}}=1$ and $\alpha_{\mathrm{forces}}=100$. An illustration of the main components of the HydraGNN architecture used for MTL is described in Figure 6.
@@ 3.5 Scalable HPO
To perform HPO at large scale, we used DeepHyper [80], an open-source Python package designed for optimizing hyperparameters, searching for optimal neural architectures. Specifically, we used asynchronous Bayesian optimization that continuously refines a surrogate model by sampling hyperparameter configurations. The efficacy of DeepHyper’s asynchronous Bayesian optimization has been demonstrated across various DL benchmarks, outperforming methods such as random search, genetic algorithms, and Hyperband in environments equipped with CPUs and GPUs. In the DeepHyper setup, a manager node refines the surrogate model and suggests promising configurations while worker nodes perform the evaluations. Our approach uses a centralized architecture with process-based parallelism, optimizing the allocation of tasks across computing nodes to avoid bottlenecks.
@@
3.6 Scalable UQ with GNN Ensembles
Ensemble methods are widely utilized in UQ to compile predictions from various models, termed ensemble members, into a unified forecast. The goal of these methods is to enhance model generalization by drawing on the diverse capabilities of each individual model [84]. To promote a varied set of predictions, practices such as different model initializations, techniques like Bagging and Boosting, and the integration of diverse network architectures are used. Research conducted by Egele et al. [85] showed that expanding the variety of network architectures within an ensemble can improve the diversity, thereby increasing the precision of uncertainty assessments. They also developed a technique for concurrently training multiple candidate models, which optimizes the use of computational resources. Ensemble methods are acknowledged for their ability to deliver reliable uncertainty estimates and their ease of implementation and scalability, making them practical for various UQ applications.
@@ 3.6 Scalable UQ with GNN Ensembles
and the uncertainty is measured as the standard deviation (STD),
@@
4 Performance measurements
4.1 Setup
For strong and weak scaling studies we utilize three different sizes of GFM architectures, denoted as SMALL, MEDIUM, and LARGE. They differ in the total number of parameters, ranging from approximately 60,000 to 163 million. Table 3 provides details about the three model sizes.
Experiments were conducted on two DOE supercomputers: Frontier at ORNL and Perlmutter at NERSC. Both systems provide state-of-the-art GPU-based heterogeneous architectures. Frontier, located at the Oak Ridge Leadership Computing Facility at ORNL, is currently the world’s fastest supercomputer [86]. It comprises a total of 9,408 compute nodes, each featuring a single 64-core AMD EPYC 7763 (Milan) CPU and four AMD Instinct MI250X GPU accelerators, effectively providing eight GPU units per node. Running with one rank per GPU unit, each rank has 64 GB of DDR4 (CPU) and 64 GB of HBM2e (GPU) memory.
@@ 4.1 Setup
To aid in monitoring HydraGNN execution in real-time for a subset of the analysis carried out on Frontier, an AMD Research utility, Omnistat [87], was used to sample a variety of GPU telemetry metrics including occupancy, high-bandwidth memory (HBM) usage, power, temperature, energy consumption, and clock/memory frequencies on a per GCD basis across all nodes assigned to an individual run. This Python-based utility was executed entirely in user-space implemented as a Prometheus client on each assigned compute node and combines low-overhead sampling via AMD’s system management interface (SMI) at fixed intervals with a temporary Prometheus server [88] instantiated on one CPU core of the master compute host per batch job. Minimal job overhead (less than 0.5%) was observed when running HydraGNN training with this approach for sampling intervals down to one second.
@@
Table 3: GNN model sizes used for strong and weak scaling tests on NERSC-Perlmutter and OLCF-Frontier.
4.2 I/O performance for reading large data
In Section 3.3, we described using the ADIOS scientific data management library for fast storage and retrieval of large training data. In this section, we show the performance of reading large data in HydraGNN for training models.
Of all the datasets used in this study, the Open Catalyst 2020 dataset is the largest in terms of number of atomistic structures, storage size of the dataset, and number of files across which data is stored. The original dataset consists of over 50,000 files. The dataset was pre-processed into ADIOS and was configured to use just over 50 ADIOS sub-files, which led to a 1000$\times$ reduction in the metadata footprint.
@@ 4.2 I/O performance for reading large data
4.3 HydraGNN Training Scaling Results
We now analyze the scaling performance of HydraGNN on Frontier and Perlmutter. We present weak and strong scaling trends, along with a breakdown of component operations in HydraGNN as we scale it up. Experiments were performed with up to 2,048 nodes on Frontier and 256 nodes on Perlmutter using the three model sizes discussed in Table 3.
4.3.1 Weak scaling
For the weak scaling runs, we configured each GPU to process 3,500 atomistic structures equally. Fig. 9 shows the weak scaling performance on Perlmutter and Frontier as we vary the number of GPUs used for the training. The reported time represents the average training time per epoch. We conducted experiments with up to 2,048 GPUs on Frontier and 1,024 GPUs on Perlmutter. The number of GPUs on Perlmutter was dictated by constraints on available node hours. We observe that the parallel efficiency of weak scaling experiments drops as we increase the number of GPUs beyond 256 for both Perlmutter and Frontier. This is attributed to increased communication costs as we scale the number of GPUs and the overhead associated with using varying graph sizes.
Figure 9: Weak scaling of HydraGNN multitasking pre-training on a problem of size 3,500 atomistic structures per GPU for (top) Frontier and (bottom) Perlmutter.
@@ 4.3.2 Strong scaling
Figure 11: Strong scaling of HydraGNN multitasking pre-training on a problem of 120 million graphs on Frontier and 2 million graphs on Perlmutter with three GNN model sizes.
The drop in scaling performance is attributed to load imbalance - an artifact of small graph sizes. As shown in Fig. 1, we use a diverse dataset where graph sizes vary by up to 400 nodes, and the number of edges in the larger graphs exceeds 12,500. This results in an imbalanced workload among GPUs in each batch, causing some GPUs to finish training before others. As GPUs must synchronize for exchanging model weights, the runtime is dominated by the GPUs that must train on larger graphs. Effectively, this leads to sub-standard utilization of compute resources and poses a challenge towards achieving high-performant, scalable training. Thereofre, while training on large volumes of data can help develop robust models because of the diverse nature of data, the computational performance may suffer as the workload can vary greatly.
@@
The EGNN model we used is particularly vulnerable to this problem. The time required for forward calculations in EGNN is directly proportional to the number of edges in the graph sample. For datasets with highly variable edge counts between graph samples, the likelihood of load imbalance between GPUs increases. Fig. 12 illustrates the time spent on the forward task in the EGNN model with different model sizes. We observe an almost linear relationship between forward time and graph size at each batch (measured by the number of edges). The SMALL models show large variances on both machines, which is expected due to system noise being more consequential for smaller model sizes. Significant performance differences (e.g., the difference between minimum and maximum time) are observed due to the varying graph sizes in our datasets. However, for other tasks (data loading and backward), we do not observe a similar correlation, as they are agnostic of the graph size. Fig. 13 illustrates the average percentage of time spent waiting during three tasks: data loading, forward pass, and backward pass. It highlights a significant waiting period during the forward pass, primarily due to varying graph sizes. This waiting time increases as the disparity in graph sizes among GPUs grows. Other tasks, such as data loading and backward pass, also involve waiting time, but to a much lesser extent.
@@
where $T_{max}$ and $T_{avg}$ represent the maximum runtime and the average runtime for training an epoch, respectively, among all computing resources (GPUs in our case). These times represent the time to perform training (forward and backward calculations) and do not include wait times during synchronization. For a well-balanced workload, LIF approaches 1.0 from above, whereas it increases as the workload imbalance increases. Fig. 14 presents the LIF scores that show the imbalance among processes. The trend remains consistent: while data loading and backward pass exhibit nearly balanced workloads (with scores close to 1.0), the forward pass shows imbalanced workload characteristics as it deviates from 1.0.